\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2} \def\path#1{#1}\fi

\bibitem{zadeh1965fuzzy}
L.~A. Zadeh, Fuzzy sets, Information and {C}ontrol 8~(3) (1965) 338--353.

\bibitem{dubois1978h}
D.~Dubois, H.~Prade, Operations on fuzzy numbers, International {J}ournal of
  {S}ystems {S}cience 9~(6) (1978) 613--626.

\bibitem{hanss2005applied}
M.~Hanss, Applied {F}uzzy {A}rithmetic: {A}n {I}ntroduction with {E}ngineering
  {A}pplications, Springer {S}cience $\with$ {B}usiness {M}edia, 2005.

\bibitem{klir1995}
G.~Klir, B.~Yuan, Fuzzy {S}ets and {F}uzzy {L}ogic: {T}heory and
  {A}pplications, Vol.~4, Prentice {H}all {N}ew {J}ersey, 1995.

\bibitem{TAsmus2017}
T.~da~Cruz~Asmus, G.~P. Dimuro, B.~R.~C. Bedregal, On two-player
  interval-valued fuzzy {B}ayesian games, International {J}ournal of
  {I}ntelligent {S}ystems 32~(6) (2017) 557--596.

\bibitem{kosinski2017}
P.~Prokopowicz, J.~M. Czerniak, D.~Mikolajewski, L.~Apiecionek, D.~Slezak
  (Eds.), Theory and Applications of Ordered Fuzzy Numbers - {A} Tribute to
  Professor Witold Kosi{\'{n}}ski, Vol. 356 of Studies in Fuzziness and Soft
  Computing, Springer, 2017.

\bibitem{Roldan-Lopez-de-Hierro18}
A.~{Rold{\'{a}}n{-}L{\'{o}}pez{-}de{-}Hierro}, C.~Rold{\'{a}}n, F.~Herrera, On
  a new methodology for ranking fuzzy numbers and its application to real
  economic data, Fuzzy Sets and Systems 353 (2018) 86--110.

\bibitem{wang2014total}
W.~Wang, Z.~Wang, Total orderings defined on the set of all fuzzy numbers,
  Fuzzy {S}ets and {S}ystems 243 (2014) 131--141.

\bibitem{zumelzu2022}
N.~Zumelzu, B.~Bedregal, E.~Mansilla, H.~Bustince, R.~Diaz, Admissible orders
  on fuzzy numbers, IEEE {T}ransactions on {F}uzzy {S}ystems 30~(11) (2022)
  4788--4799.

\bibitem{bustince2013generation}
H.~Bustince, J.~Fern{\'a}ndez, A.~Koles{\'a}rov{\'a}, R.~Mesiar, Generation of
  linear orders for intervals by means of aggregation functions, Fuzzy Sets and
  Systems 220 (2013) 69--77.

\bibitem{Annax21}
A.~A. {De Lima}, E.~S. Palmeira, B.~Bedregal, H.~Bustince, Multidimensional
  fuzzy sets, IEEE Transactions on Fuzzy Systems 29~(8) (2021) 2195--2208.

\bibitem{Laura16b}
L.~{De Miguel}, H.~Bustince, B.~Pekala, U.~Bentkowska, I.~A. {da Silva},
  B.~R.~C. Bedregal, R.~Mesiar, G.~Ochoa, Interval-valued {A}tanassov
  intuitionistic {OWA} aggregations using admissible linear orders and their
  application to decision making, {IEEE} Transaction on Fuzzy Systems 24~(6)
  (2016) 1586--1597.

\bibitem{matzenauer2021strategies}
M.~Matzenauer, R.~Reiser, H.~Santos, B.~Bedregal, H.~Bustince, Strategies on
  admissible total orders over typical hesitant fuzzy implications applied to
  decision making problems, International {J}ournal of {I}ntelligent {S}ystems
  36~(5) (2021) 2144--2182.

\bibitem{hamilton1853wr}
W.~R. Hamilton, Lectures on Quaternions: Containing a Systematic Statement of a
  New Mathematical Method; of which the Principles Were Communicated in 1843 to
  the Royal Irish Academy; and which Has Since Formed the Subject of Successive
  Courses of Lectures, Delivered in 1848 and Subsequent Years in the Halls of
  Trinity College, Dublin: With Numerous Illustrative Diagrams, and with Some
  Geometrical and Physical Applications, Hodges and Smith, 1853.

\bibitem{cayley1858ii}
A.~Cayley, Ii. a memoir on the theory of matrices, Philosophical {T}ransactions
  of the Royal {S}ociety of London~(148) (1858) 17--37.

\bibitem{grassmann1844lineale}
H.~Grassmann, Die lineale Ausdehnungslehre ein neuer Zweig der Mathematik:
  dargestellt und durch Anwendungen auf die {\"u}brigen Zweige der Mathematik,
  wie auch auf die Statik, Mechanik, die Lehre vom Magnetismus und die
  Krystallonomie {E}rl{\"a}utert, Vol.~1, O. Wigand, 1844.

\bibitem{peano1887integrazione}
G.~Peano, Integrazione per serie delle equazioni differenziali lineari, Atti
  Accad. Sci. Torino 22 (1887) 437--446.

\bibitem{colloc2020fvsoomm}
J.~Colloc, Fvsoomm a fuzzy vectorial space model and method of personality,
  cognitive dissonance and emotion in decision making, Information 11~(4)
  (2020) 229.

\bibitem{colloc2017epice}
J.~Colloc, R.~A. Yameogo, P.~F. Summons, Y.~Shen, M.~Park, J.~E. Aronson, Epice
  an emotion fuzzy vectorial space for time modeling in medical decision, in:
  Proceedings of the 1st International Conference on Internet of Things and
  Machine Learning, 2017, pp. 1--7.

\bibitem{katsaras1977fuzzy}
A.~Katsaras, D.~B. Liu, Fuzzy vector spaces and fuzzy topological vector
  spaces, Journal of Mathematical Analysis and Applications 58~(1) (1977)
  135--146.

\bibitem{cristescu1976ordered}
R.~Cristescu, Ordered vector spaces and linear operators, Routledge, 1976.

\bibitem{milfont2021aggregation}
T.~Milfont, I.~Mezzomo, B.~Bedregal, E.~Mansilla, H.~Bustince, Aggregation
  functions on $n$-dimensional ordered vectors equipped with an admissible
  order and an application in multi-criteria group decision-making,
  International {J}ournal of {A}pproximate {R}easoning 137 (2021) 34--50.

\bibitem{costa2015generalized}
T.~Costa, Y.~Chalco-Cano, W.~A. Lodwick, G.~N. Silva, Generalized interval
  vector spaces and interval optimization, Information Sciences 311 (2015)
  74--85.

\bibitem{vasantha1993semivector}
W.~Vasantha~Kandasamy, Semivector spaces over semifields, Zeszyty Nauwoke
  Politechniki 17 (1993) 43--51.

\bibitem{kandasamy2002smarandache}
W.~V. Kandasamy, Smarandache Semirings, Semifields, And Semivector Spaces,
  Infinite Study, 2002.

\bibitem{rodriguez2016position}
R.~M. Rodr{\'\i}guez, B.~Bedregal, H.~Bustince, Y.~Dong, B.~Farhadinia,
  C.~Kahraman, L.~Mart{\'\i}nez, V.~Torra, Y.~Xu, Z.~Xu, F.~{H}errera, A
  position and perspective analysis of hesitant fuzzy sets on information
  fusion in decision making. towards high quality progress, Information Fusion
  29 (2016) 89--97.

\bibitem{beliakov2016practical}
G.~Beliakov, H.~B. Sola, T.~C. S{\'a}nchez, A practical guide to averaging
  functions, Springer, 2016.

\bibitem{dimuro2017new}
G.~P. Dimuro, H.~Bustince, J.~Fernandez, R.~Mesiar, B.~Bedregal, New results on
  pre-aggregation functions: introducing (light) pre-t-norms, in: 2017 Joint
  17th World Congress of International Fuzzy Systems Association and 9th
  International Conference on Soft Computing and Intelligent Systems
  (IFSA-SCIS), IEEE, 2017, pp. 1--6.

\bibitem{yang2020some}
R.~Yang, L.~Jin, D.~Paternain, R.~R. Yager, R.~Mesiar, H.~Bustince, Some
  preference involved aggregation models for basic uncertain information using
  uncertainty transformation, Journal of Intelligent \& Fuzzy Systems 39~(1)
  (2020) 325--332.

\bibitem{da2022towards}
T.~da~Cruz~Asmus, G.~P. Dimuro, B.~Bedregal, J.~A. Sanz, R.~Mesiar,
  H.~Bustince, Towards interval uncertainty propagation control in bivariate
  aggregation processes and the introduction of width-limited interval-valued
  overlap functions, Fuzzy Sets and Systems 441 (2022) 130--168.

\bibitem{Reiser13}
R.~H.~S. Reiser, B.~R.~C. Bedregal, M.~Baczynski, Aggregating fuzzy
  implications, Information Sciences 253 (2013) 126--146.

\bibitem{de2021affine}
A.~F. {Rold{\'a}n-L{\'o}pez-de-Hierro}, C.~Rold{\'a}n, H.~Bustince,
  J.~Fern{\'a}ndez, I.~Rodriguez, H.~Fardoun, J.~Lafuente, Affine construction
  methodology of aggregation functions, Fuzzy Sets and Systems 414 (2021)
  146--164.

\bibitem{dimuro2020state}
G.~P. Dimuro, J.~Fern{\'a}ndez, B.~Bedregal, R.~Mesiar, J.~A. Sanz, G.~Lucca,
  H.~Bustince, The state-of-art of the generalizations of the {C}hoquet
  integral: from aggregation and pre-aggregation to ordered directionally
  monotone functions, Information Fusion 57 (2020) 27--43.

\bibitem{galar2013aggregation}
M.~Galar, A.~Jurio, C.~Lopez-Molina, D.~Paternain, J.~Sanz, H.~Bustince,
  Aggregation functions to combine {RGB} color channels in stereo matching,
  Optics Express 21~(1) (2013) 1247--1257.

\bibitem{gonzalez2014use}
L.~Gonz{\'a}lez-Jaime, M.~Nachtegael, E.~Kerre, H.~Bustince, Use of idempotent
  functions in the aggregation of different filters for noise removal, in:
  Knowledge Engineering and Management, Springer, 2014, pp. 495--507.

\bibitem{OliveiraADYRB22}
L.~de~Souza~Oliveira, A.~Argou, R.~Dilli, A.~C. Yamin, R.~Reiser, B.~R.~C.
  Bedregal, Exploring fuzzy set consensus analysis in iot resource ranking,
  Engineering Applications of Artificial Intelligence 109 (2022) 104617.

\bibitem{Batista22}
T.~V.~V. Batista, B.~R.~C. Bedregal, R.~M. Moraes, Constructing multi-layer
  classifier ensembles using the {C}hoquet integral based on overlap and
  quasi-overlap functions, Neurocomputing 500 (2022) 413--421.

\bibitem{Costa18}
V.~S. Costa, A.~D.~S. Farias, B.~R.~C. Bedregal, R.~H.~N. Santiago, A.~M.
  de~Paula~Canuto, Combining multiple algorithms in classifier ensembles using
  generalized mixture functions, Neurocomputing 313 (2018) 402--414.

\bibitem{nolasco2019wavelet}
D.~H. Nolasco, F.~B. Costa, E.~S. Palmeira, D.~K. Alves, B.~Bedregal, T.~O.
  Rocha, R.~L. Ribeiro, J.~C. Silva, Wavelet-fuzzy power quality diagnosis
  system with inference method based on overlap functions: {C}ase study in an
  {A}{C} microgrid, Engineering {A}pplications of {A}rtificial {I}ntelligence
  85 (2019) 284--294.

\bibitem{fumanal2021motor}
J.~Fumanal-Idocin, Y.-K. Wang, C.-T. Lin, J.~Fern{\'a}ndez, J.~A. Sanz,
  H.~Bustince, Motor-imagery-based brain-computer interface using signal
  derivation and aggregation functions, IEEE Transactions on Cybernetics
  (2021).

\bibitem{saleh2018learning}
E.~Saleh, A.~Valls, A.~Moreno, P.~Romero-Aroca, V.~Torra, H.~Bustince, Learning
  fuzzy measures for aggregation in fuzzy rule-based models, in: International
  Conference on Modeling Decisions for Artificial Intelligence, Springer, 2018,
  pp. 114--127.

\bibitem{BustinceMFGPADB21}
H.~Bustince, R.~Mesiar, J.~Fern{\'{a}}ndez, M.~Galar, D.~Paternain, A.~H.
  Altalhi, G.~P. Dimuro, B.~R.~C. Bedregal, Z.~Tak{\'{a}}c, d-{C}hoquet
  integrals: {C}hoquet integrals based on dissimilarities, Fuzzy Sets and
  Systems 414 (2021) 1--27.

\bibitem{LuccaSDBB20}
G.~Lucca, J.~A. Sanz, G.~P. Dimuro, B.~R.~C. Bedregal, H.~Bustince, A proposal
  for tuning the {\(\alpha\)} parameter in {C}\({}_{\mbox{{\(\alpha\)}
  }}\){C}-integrals for application in fuzzy rule-based classification systems,
  Natural Computing 19~(3) (2020) 533--546.

\bibitem{LuccaSDBMKB16}
G.~Lucca, J.~A. Sanz, G.~P. Dimuro, B.~R.~C. Bedregal, R.~Mesiar,
  A.~Koles{\'{a}}rov{\'{a}}, H.~Bustince, Preaggregation functions:
  Construction and an application, {IEEE} Transaction on Fuzzy Systems 24~(2)
  (2016) 260--272.

\bibitem{zhu2019nested}
C.~Zhu, L.~Jin, R.~Mesiar, R.~R. Yager, D.~Paternain, H.~Bustince, Nested
  formulation paradigms for induced ordered weighted averaging aggregation for
  decision-making and evaluation, International Journal of Intelligent Systems
  34~(11) (2019) 3046--3057.

\bibitem{apostol2007calculus}
T.~M. Apostol, Calculus. Volumen 1 y 2, 2007.

\bibitem{rudin1964principles}
W.~Rudin, Principles of {M}athematical {A}nalysis, Vol.~3, McGraw-{H}ill {N}ew
  {Y}ork, 1964.

\bibitem{valentina2021}
V.~Rosas, B.~Bedregal, J.~Canumán, R.~Díaz, E.~Mansilla, N.~Zumelzu, A new
  total order for triangular fuzzy numbers with an application, in: Anais do
  {V}{I} {W}orkshop-{E}scola de {I}nform\'atica {T}e\'orica, SBC, Porto
  {A}legre, RS, {B}rasil, 2021, pp. 17--24.
\newblock \href {https://doi.org/10.5753/weit.2021.18917}
  {\path{doi:10.5753/weit.2021.18917}}.

\end{thebibliography}

\documentclass[preprint,12pt]{elsarticle}
\journal{xxxxxxxx}
\usepackage{anyfontsize,t1enc}
\usepackage{lscape}
\usepackage{array}
\usepackage{booktabs}

\usepackage[noend]{algpseudocode}
\usepackage{here}
\let\oldReturn\Return
\renewcommand{\Return}{\State\oldReturn}

\usepackage{lineno} % to enumerate the lines

\usepackage{babelbib,multicol}

\usepackage[caption=false,font=normalsize,labelfon
t=sf,textfont=sf]{subfig}

\usepackage[english]{babel}

\usepackage{amsmath,dsfont,amsthm,cancel,amsfonts,amssymb,mathrsfs,amsthm,cmll,nicefrac}
\usepackage{textcomp}

\usepackage{algorithm}

\usepackage{color}

\usepackage{enumitem}
\usepackage{graphics}
\usepackage{lineno}

\usepackage{hyperref}
\usepackage{MnSymbol,stmaryrd} %,yhmath} 
\usepackage{manfnt}

\usepackage{tikz}
\usetikzlibrary{automata} 

\definecolor{rojo}{RGB}{172,6,6}
\definecolor{negro}{RGB}{25,6,6}

\newcolumntype{M}[1]{>{\hbox to #1\bgroup\hss$}l<{$\egroup}}

\makeatletter
\newcommand\@brcolwidth{2.9em}
\newenvironment{brmatrix}{%
    \left(%
    \hskip-\arraycolsep
    \new@ifnextchar[\@brarray{\@brarray[\@brcolwidth]}%
}{%
    \endarray
    \hskip -\arraycolsep
    \right)%
}
\def\@brarray[#1]{\array{r*\c@MaxMatrixCols {M{#1}}}}
\makeatother
\DeclareMathOperator{\Maxx}{Max}
\DeclareMathOperator{\Minn}{Min}
\DeclareMathOperator{\supp}{supp}

\DeclareMathOperator{\TFN}{TFN}
\DeclareMathOperator{\OT}{OT}
\providecommand{\leqot}[0]{\leq_{\OT}}
\providecommand{\geqot}[0]{\geq_{\OT}}
\newcommand{\ITEM}[1]{\noindent{#1}}
\providecommand{\Vector}[1]{\overrightarrow{#1}}
\providecommand{\NFT}[1]{(a_{#1},b_{#1},c_{#1})}

\providecommand{\crispy}[1]{\widetilde{#1}}

\providecommand{\cortea}[1]{\nicefrac{\Large{#1}}{\alpha}}

\providecommand{\operador}[3]{#1:#2\longrightarrow #3}
\providecommand{\con}[1]{\mathds{#1}}

\providecommand{\Sup}[1]{\underset{#1}{\sup~}}
\providecommand{\NFR}[0]{\mathcal{F}(\mathds{R})}

\providecommand{\NFRN}[1]{\mathcal{F}_{-}(\mathds{R})}

\providecommand{\TFNR}{\mathfrak{T}(\con{R})}

\providecommand{\suma}[2]{\overset{#1}{\underset{#2}{\sum}}}

\newcommand{\prt}[1]{\langle #1\rangle}

\newtheorem{teorema}{Theorem}[section]
\newtheorem*{tricotomia}{$OT$-Trichotomy law for $\TFNR$}
\newtheorem{proposicion}{Proposition}[section]
\newtheorem{corolario}{Corollary}[section]

\newtheorem{definicion}{Definition}[section]
\newtheorem{observacion}{Remark}[section]

\newtheorem{ejemplo}{Example}[section]

\begin{document}

\begin{frontmatter}

\title{Averaging functions on triangular fuzzy numbers}

%% Group authors per affiliation:
\author{Roberto D\'iaz$^a$, Aldryn Aparcana$^{b}$, Diego Soto$^c$, Nicolás Zumelzu$^c$, Jos\'e Canum\'an$^{d}$, Alvaro Mella$^{c}$, Edmundo Mansilla$^{c}$ and Benjam\'in Bedregal$^e$}

\cortext[]{Corresponding author: nicolas.zumelzu@umag.cl }

\ead{roberto.diaz@ulagos.cl, aldryn.aparcana@unica.edu.pe,disoto@umag.cl, nicolas.zumelzu@umag.cl, jose.canuman@umag.cl, alvaro.mella@umag.cl, edmundo.mansilla@umag.cl and bedregal@dimap.ufrn.br}

\address[a]{Departamento de Ciencias Exactas, Universidad de Los Lagos, Osorno, Chile}

\address[b]{Facultad de Ciencias, Universidad Nacional San Luis Gonzaga, Ica, Peru}

\address[c]{Departamento de Matem\'atica y F\'isica, Universidad de Magallanes, Punta Arenas, Chile}

\address[d]{Departamento de Ingenier\'ia en Computaci\'on, Universidad de Magallanes, Punta Arenas, Chile}
\address[e]{Departamento de Inform\'{a}tica e Matem\'{a}tica Aplicada, Universidade Federal do Rio Grande do Norte, Natal, Brazil}

\begin{abstract} Admissible orders on fuzzy numbers are total orders which refine a basic and well-known partial order on fuzzy numbers. In this work, we define an admissible order on triangular fuzzy numbers (i.e. $\TFN$'s) and study some fundamental properties with its arithmetic and their relation with this admissible order. In addition, we also introduce the concepts of average function on $\TFN$'s and studies a generalized structure of the vector spaces.  In particular we consider the case of $\TFN$'s.

\begin{keyword}
Triangular fuzzy numbers \sep Orders on fuzzy numbers \sep admissible orders
 \sep vector space.
\end{keyword}

\end{abstract}

\end{frontmatter}

\section{Introduction}

Zadeh in \cite{zadeh1965fuzzy} proposed the concept of fuzzy sets to formalize and model the ambiguities and inaccuracies of language such as temperature, age, and speed. A class of fuzzy sets, called of fuzzy numbers, were introduced in \cite{zadeh1965fuzzy} to model a quantity that is imprecise, rather than exact, as it is in all traditional mathematics. This structure has been studied in countless works, and has been dedicated to: arithmetic operations  and other related notions for fuzzy numbers and their properties (see \cite{dubois1978h,hanss2005applied,klir1995}). Additionally, there are several proposal of partial orders (totals and non totals) for fuzzy numbers or a subclass of the fuzzy numbers as for example in \cite{TAsmus2017,kosinski2017,Roldan-Lopez-de-Hierro18,wang2014total}. More recently, Zumelzu et al. in \cite{zumelzu2022}, in the light of the notion of admissible orders in the context of extensions of fuzzy set theory as in \cite{bustince2013generation,Annax21,Laura16b,matzenauer2021strategies},  proposed the notion of admissible orders for fuzzy numbers as total orders which refines the order of Klir and Yuan in \cite{klir1995}.

To talk about vector spaces, we have to highlight the hard work of mathematicians who allowed the definition and axiomatization of this concept. Among them, we can find \cite{hamilton1853wr}, that by his definition of the complex numbers and the operations of the quaternions, one of the main elements of the vector space could be defined i.e. the geometric vectors, where later on \cite{cayley1858ii}. As is known, the theory of matrices was postulated from systems of linear equations, the definition of the matrix and its properties, within these definitions, implicitly, provides an example of vector space. In \cite{grassmann1844lineale} different crucial concepts are developed in vector spaces, such as linear dependence, linear independence, among others, also defining the concept of vector in terms of vector spaces without formalizing the latter, so that in \cite{peano1887integrazione} the formal definition of vector space is presented in an axiomatic way.

There are several extensions of vector spaces in fuzzy logic called fuzzy vector spaces \cite{colloc2020fvsoomm,colloc2017epice,katsaras1977fuzzy}. Regarding its applications we can find, for example, application to cognitive dissonance and decision making, Time fuzzy vector space (i.e. TFVS) modeling of emotion in the purchase decision, application to gaming addiction, modeling of time in medical decisions \cite{colloc2020fvsoomm,colloc2017epice}. Regarding some generalizations of ordered vector spaces \cite{cristescu1976ordered} we can find that of ordered semi-vector spaces over a weak semi-field \cite{milfont2021aggregation}, $n$-dimensional fuzzy sets over a non-empty set, among others \cite{milfont2021aggregation, costa2015generalized,vasantha1993semivector, kandasamy2002smarandache,rodriguez2016position}.

Namely, the aggregation functions play an important role in several areas, including fuzzy logic, decision-making, expert systems, risk analysis, and image processing.  The aggregation functions combine input values into a single output value, which represents all the inputs. Typical examples are weighted means, medians, OWA functions, $t$-norms and $t$-conorms. But there are many other aggregation functions and infinitely many members in most families. In general, there are several contributions to the study of aggregation functions, namely { \cite{beliakov2016practical, dimuro2017new,   yang2020some}, which, we can find them in multi-criteria decision making problems \cite{beliakov2016practical, da2022towards}, connectives in fuzzy logic \cite{beliakov2016practical,Reiser13},  image processing \cite{de2021affine, dimuro2020state,  galar2013aggregation, gonzalez2014use}, in IoT \cite{OliveiraADYRB22}, classifier ensemble \cite{Batista22,Costa18}, forest fire detection \cite{da2022towards}, power quality diagnosis \cite{nolasco2019wavelet}, motor-imagery-based brain-computer interface \cite{fumanal2021motor}, in medicine to estimate the risks of a person to develop a disease \cite{saleh2018learning}, the increasing interest in the study of this topic by defining new classes of aggregation functions \cite{de2021affine}. Among some of the classes to be defined we can find, for example, the pre-aggregation, variants of the Choquet integral and OWA functions \cite{dimuro2020state,BustinceMFGPADB21,LuccaSDBB20,LuccaSDBMKB16,zhu2019nested}.}

In this work, we extend the theory of functions of average type. The space studied here is that given by the cartesian product of fuzzy triangular numbers denoted by  $\TFNR^n$, considering as a set of scalars the field of crisp fuzzy numbers denoted by $\crispy{\con{R}}$. We prove that this set fulfills properties such as associativity, commutativity, additive neutral both for the addition of elements of $\TFNR^n$ as by the product of scalars with elements of $\TFNR^n$. We introduce the concept of increasing average type function proving properties and showing some examples. To study these increasing functions we have defined a linear order on the fuzzy triangular numbers by studying arithmetic properties related to the order. 

This paper is organized as follows: In Section \ref{preliminares}, in addition to establishing the notation used, we recall some essential notions for the remaining sections. In Section \ref{totalorder} study of arithmetic properties and order. We introduce the ordered vector spaces on $\TFN$ in Section \ref{TriangularfuzzySpace}. Section \ref{Averaging-Functions}, we define aggregation functions of the average type on $\TFN$. In addition, we give examples and a characterization of it. 

\section{Preliminaries}\label{preliminares}

In this section, we provide the fuzzy number context and the main {concepts and results which are necessaries in  the remain of this paper}. Let $\con{R}^3 = \con{R}\times\con{R}\times\con{R}$, with $\con{R}$ the set of real numbers, $\con{Q}$ and $\con{I}$ denote the set of rational and irrational numbers, respectively. The notation $<$, $>$, $\leq$ or $\geq$ stands for usual order of $\con{R}$, and $[a,b]$, $]a,b]$, $[a,b[$ and $]a,b[$ the intervals of $\con{R}$ closed, right-closed, right-open,  and open respectively (see \cite{apostol2007calculus,
rudin1964principles}). 

{
\begin{definicion} \cite{klir1995}
 A fuzzy set $A$ over a reference non empty set $X$ is a function $A:X\rightarrow [0,1]$.
\end{definicion}
}

%\vspace{5mm}
\begin{definicion}\cite{klir1995} \label{def-FN} A fuzzy set $A$ over $\con{R}$ is called a fuzzy number if it satisfies the following conditions
\begin{enumerate}
\item[(i)] $A$ is normal, i.e., $\Sup{x\in\con{R}} A(x)=1$;
\item[(ii)] $\cortea{A}$ is a closed interval for every $\alpha\in ~]0, 1]$;
\item[(iii)] $\supp A$, i.e., the support of $A$, is bounded,
\end{enumerate}
where $\nicefrac{A}{\alpha}$ is the $\alpha$-level (or $\alpha$-cut) of $A$.
\end{definicion}

A fuzzy number $A$ is a crisp fuzzy number or in short crisp if there exists $r\in\mathds{R}$ such that $A(r)=1$ and $A(x)=0$ for each $x\neq r$. In this case we will denote $A$ by $\crispy{r}$. Finally, $\NFR$ will denote the set of all fuzzy numbers and $\crispy{\con{R}}$, $\crispy{\con{Q}}$ and $\crispy{\con{I}}$ will denote the set for all crisp fuzzy number{s} with $r\in\con{R}$, $r\in\con{Q}$ and  $r\in\con{I}${, respectively}. \vspace{5mm}

\begin{proposicion}\cite{dubois1978h} Let $A,B\in \mathcal{F}(\con{R})$ then the fuzzy sets $A+B$, $A-B$, $A\cdot B$ and $A\div B$ defined {for each $x\in \con{R}$} by
\begin{enumerate}
\item $(A+B)(x)=\Sup{x=y+z} \min\{A(y),B(z)\}$,
\item $(A-B)(x)=\Sup{x=y-z} \min\{A(y),B(z)\}$,
\item $(A\cdot B)(x)=\Sup{x=yz} \min\{A(y),B(z)\}$,
\item $(A\div B)(x)=\Sup{x=\frac{y}{z}} \min\{A(y),B(z)\},$ since $0\not\in\supp B$,
\end{enumerate}
 is a fuzzy number.
\end{proposicion}
\vspace{5mm}

%\vspace{5mm}
\begin{definicion}\cite{klir1995} \label{deftriangularnumero}
A fuzzy number $A$, is called a triangular fuzzy number, in short $\TFN$, if there is $(a,b,c)\in \mathds{R}^3$ such that $a\leq b\leq c$ and
\begin{equation*}
A(x)=\left\{\begin{array}{ll}
1, & \textrm{ if }x=b,\\
\frac{x-a}{b-a}, & \textrm{ if } x \in~ ]a,b[,\\
\frac{c-x}{c-b}, & \textrm{ if }x \in~]b,c[,\\
0,& \textrm{ other cases. }
\end{array}\right.
\end{equation*}
\end{definicion}
%\vspace{5mm}
In this work, we represent the fuzzy triangular numbers by the triple $(a,b,c)$ and denote the set of all triangular fuzzy numbers by  $\TFNR$, namely $$\TFNR=\{A\in\NFR:A\textrm{ is a triangular fuzzy numbers}\}.$$ %(-a,0,a)
{Observe that each triple $(a,b,c)\in \mathds{R}^3$ satisfying $a\leq b\leq c$ denotes a $\TFN$.}
\begin{proposicion}\cite{dubois1978h}\label{pro-oper}%\cite{Klir1995}
If $A=(a_1,b_1,c_1)$ and $B=(a_2,b_2,c_2)$ be are elements in $\TFNR$ and $\crispy{r}\in\crispy{\con{R}}$.  Then
\begin{enumerate}
\item $A+B=(a_1+a_2,b_1+b_2,c_1+c_2)$,
\item $A-B=(a_1-c_2,b_1-b_2,c_1-a_2)$,
\item $\crispy{r}A=\begin{cases}(ra_1,rb_1,rc_1)\textrm{ , if }r\geq 0,\\
(rc_1,rb_1,ra_1) \textrm{ , if }r<0.\end{cases}$
\end{enumerate}
i.e. the set {$\TFNR$ is} closed under addition, substraction and scalar product. Note that $(-1)A=-A=-(a_1,a_2,a_3)=(-a_3,-a_2,-a_1)$.
\end{proposicion}
\begin{corolario}
Let $A$ be a fuzzy {number}. Then $\underset{n-\textrm{times}}{\underbrace{A+A+...+A}}=\crispy{n}A$.
\end{corolario}
\begin{observacion}
The product and division, i.e.,  $A\cdot B$ and $A\div B$,  of triangular fuzzy numbers only are $\TFN$'s if and only if $A$ or $B$ is {a crisp} fuzzy number (obviously $0\not\in\supp{B}$ in the case of division).
\end{observacion}

 \begin{observacion}\label{rem-ult-sec2}
The addition is commutative and associative in fuzzy triangular numbers. Also, addition distributes over the product and vice versa, where the product is given in Proposition \ref{pro-oper}. Note also that the crisp fuzzy numbers with the operations of addition and multiplication are isomorphic to the field of real numbers and are therefore also a field.
\end{observacion}

\begin{definicion}\label{definicionkytfn} Let $A=(a_1,b_1,c_1)$ and $B=(a_2,b_2,c_2)$ be two $\TFN$'s. We define the relation $\leq_N$ by
$$A\leq_N B \textrm{ if and only if } a_1\leq a_2\textrm{ and }b_1\leq b_2 \textrm{ and } c_1\leq c_2.$$
We denote $A<_N B$ when $A\neq B$, i.e. $a_1\neq a_2$ or $b_1\neq b_2$ or $c_1\neq c_2$.
\end{definicion}

The Definition \ref{definicionkytfn} generalizes the usual order of real numbers, in the sense that both agree when to restrict to crisp fuzzy numbers. It is worth noting that this relation is the restriction to the $\TFN$'s of the Klir and Yuan partial order on the fuzzy numbers. In the following, we adapt the definition of admissible orders on fuzzy numbers introduced in \cite{zumelzu2022} to the context of $\TFN$'s.

\begin{definicion}\label{def-ordenadmisible}
Let $\leq_N$ and $\preceq$ be two orders on the $\TFN$'s. The order $\preceq$ is called an admissible order w.r.t. {$\prt{{\TFNR},\leq_N}$}, if
\begin{enumerate}[labelindent=\parindent, leftmargin=*,label=]
\item[(i)] $\preceq$ is a linear order on ${\TFNR}$;
\item[(ii)] for all $A$, $B$ in ${\TFNR}$, $A\preceq B$ whenever $A\leq_N B$.
\end{enumerate}
\end{definicion}

\section{Admissible order}\label{totalorder}
In this section, we focus on proving {an admissible order on $\TFNR$,} we also study properties that such linear {order} verifies relates to the addition and subtraction operations.
 \vspace{5mm}
 
\begin{definicion}\label{defOT}
  Let $(a_1,b_1,c_1)$ and $(a_2,b_2,c_2)$ be two $\TFN$'s. We define the relation $\leqot$ by
\begin{multline*}
(a_1,b_1,c_1)\leqot (a_2,b_2,c_2) \textrm{ if and only if }  \begin{cases} b_1<b_2,
\textrm{ or }\\
 b_1=b_2\textrm{ and }c_1<c_2, \textrm{ or }\\
 b_1=b_2\textrm{ and }c_1=c_2\textrm{ and }a_1\leq a_2.\end{cases}
\end{multline*}
We denote $A<_{\OT} B$ when {$A\leqot B$ and $A\neq B$, i.e.} $a_1\neq a_2$ or $b_1\neq b_2$ or $c_1\neq c_2$.
\end{definicion}

\begin{proposicion}\label{OTordentotal}
The relation $\leqot$ is an admissible order on  $\TFNR$.
\end{proposicion}

\begin{proof}
Let $A=(a_1,b_1,c_1)$, $B=(a_2,b_2,c_2)$, $C=(a_3,b_3,c_3)\in\TFNR$. 

\ITEM{1.} Reflexivity: Straightforward from Definition \ref{defOT}.

\ITEM{2.} Antisymmetry: Let $A\leqot B$ and $B\leqot A$. If $b_1\neq b_2$ then either $A<_{OT}$ B or $B<_{OT}$ A but not both. If $b_1=b_2$ and $c_1\neq c_2$ then, $A<_{OT} B$ or $B<_{OT} A$ but not both. If $b_1=b_2$ and $c_1=c_2$ and $a_1\neq a_2$ then either $A<_{OT} B$ or $B<_{OT} A$ but not both. Thereby, $a_1=a_2$, $b_1=b_2$ and $c_1=c_2$.

\ITEM{3.} {Transitivity:} If $A\leqot B$ and $B\leqot C$ then
\begin{enumerate}
\item $b_1<b_2$ or ($b_1=b_2$ and $c_1<c_2$) or ($b_1=b_2$ and $c_1=c_2$ and $a_1\leq a_2$),
\item $b_2<b_3$ or ($b_2=b_3$ and $c_2<c_3$) or ($b_2=b_3$ and $c_2=c_3$ and $a_2\leq a_3$).
\end{enumerate}
This proof consists of 9 steps, although here we will demonstrate 4.  Step 1: If $b_1<b_2$ then $b_1<b_3$ and therefore $(a_1,b_1,c_1)<_{\OT}(a_3,b_3,c_3)$. { Step 2: If $b_1<b_2$ and $b_2=b_3$ and $c_2<c_3$ then $b_1<b_2=b_3\leq c_2<c_3$, and therefore $(a_1,b_1,c_1){\leqot} (a_3,b_3,c_3)$. Step 3: If $b_1<b_2$ and $b_2=b_3$ and $c_2=c_3$ and $a_2\leq a_3$ then $b_1<b_3$. Step 4: If $b_1=b_2$ and $c_1=c_2$ and $a_1\leq a_2$ and $b_2=b_3$ and $c_2<c_3$ then $a_1\leq a_2\leq b_1=b_2=b_3\leq c_1=c_2<c_3$. Hence, $c_1<c_3$}. Therefore $(a_1,b_1,c_1)\leq_{OT}(a_3,b_3,c_3)$. 

\ITEM{4.} Totality: Let $(a_1,b_1,c_1)\neq (a_2,b_2,c_2)$. Then $a_1\neq a_2$ or $b_1\neq b_2$ or $c_1\neq c_2$. So, by Definition \ref{defOT} and trichotomy property of real numbers where for any two real numbers, $x$ and $y$, $x<y$ or $y> x$ or $x=y$, it is $(a_1,b_1,c_1)<_{OT}(a_2,b_2,c_2)$ or $(a_2,b_2,c_2)<_{OT}(a_1,b_1,c_1)$.

\ITEM{5.} {Refinement}: Let $(a_1,b_1,c_1)\leq_N (a_2,b_2,c_2)$. If $A=B$ then, since $\leqot$ is reflexive, we have that $(a_1,b_1,c_1)\leqot (a_2,b_2,c_2)$.
If $(a_1,b_1,c_1)\leq_N (a_2,b_2,c_2)$ then by Definition \ref{definicionkytfn} we have $a_1\leq_N a_2$ and $b_1\leq_N b_2$ and $c_1\leq_N c_2$ and ($a_1\neq a_2$ or $b_1\neq b_2$ or $c_1\neq c_2$). Then, if $b_1\neq b_2$ hence $b_1 <b_2$, but, if $b_1=b_2$ and $c_1\neq c_2$ then $b_1=b_2$ and $c_1< c_2$. Finally, if $b_1=b_2$ and $c_1= c_2$ and $a_1\neq a_2$ then $b_1=b_2$ and $c_1= c_2$ and $a_1< a_2$.  Thereby, $A\leqot B$.

In this way the Proposition is proven. 
\end{proof}

\begin{definicion}
Let $A\in \TFNR$. Then,
\begin{enumerate}
\item $A$ is $OT$-positive if $A >_{OT}\crispy{0}$.
\item $A$ is $OT$-negative if $A <_{OT}\crispy{0}$.
\end{enumerate}
\end{definicion}

In $\TFNR$ the elements of the form $(-a,0,a)$ with $a\leq 0$ we calling $0$-\textit{isosceles triangular {fuzzy numbers}} and  {denoted by $\TFNR_0$ to the set of all $0$-isosceles $\TFN$'s. In} particular $\crispy{0}=(0,0,0)$ belongs to $\TFNR_0$.
\begin{tricotomia}\label{positovoornegativo}
If $A\in \TFNR$ then one and only one of the following statements is true: 
\begin{enumerate}
\item $A$ is $OT$-positive;
\item $A$ is $OT$-negative;
\item $A=\crispy{0}$.
\end{enumerate}
\end{tricotomia}

\begin{proof} Straightforwardly, because
$\leqot$ is a linear order on
$\TFN$.
\end{proof}

{Observe that the trichotomy law non holds case consider the natural order  instead of $\leq_{OT}$.}

\begin{proposicion}\label{proposiciondeB}
Let $(a,b,c)$ be a $\TFN$. Then, $(a,b,c)$ is $OT$-positive if and only if $b>0$ or $(0=b< c)$. In the same way, $(a,b,c)$ is $OT$-negative if and only if $b< 0$ or $a<b=c=0$.
\end{proposicion}
\begin{proof} Straightforwardly from above definition.
\end{proof}
\begin{corolario}\label{positivonegativo}
Let $A\in\TFNR$. If $A$ is $OT$-negative then $-A$ is $OT$-positive. We note that the converse in general is not true.
\end{corolario}

\begin{ejemplo}
Let $(-1,0,1)$ is $OT$-positive. Note that $-(-1,0,1)=(-1,0,1)$ is also $OT$-positive.
\end{ejemplo}

\begin{corolario}\label{corolarioceroespecial}
Let $A\in\TFNR_0$. If $A\neq\crispy{0}$ then ${-A}$ is $OT$-positive.
\end{corolario}

\begin{proposicion}\label{proposicionrestadacasicero}
If $A\in \TFNR$ then $A-A\in \TFNR_0$.
\end{proposicion}
\begin{proof}
Let $A=(a,b,c)\in{\TFNR}$. Then $(a,b,c)-(a,b,c)=(a-c,0,c-a)$ i.e. $c-a=-a+c=-(a-c)$ but $a<c$ hence $c-a>0$. In this way the Proposition is prove.
\end{proof}
In the following corollary, we obtain a property of over-additive inverse.
\begin{corolario}\label{teoinverso}
Let $A$ be a ${\TFN}$. Then $A-A\geq_{OT}\crispy{0}$. In addition, $A-A=\crispy{0}$ if and only if $A$ is crisp.
\end{corolario}

\begin{proposicion} 
 Let $A$, $B$ two $\TFNR$. $A\leqot B$ if and only if $B-A\geqot\crispy{0}$.
\end{proposicion}

\begin{proof}
Let $\NFT{1}$, $\NFT{2}\in\TFNR$ such that $\NFT{1}\leqot\NFT{2}$. If $b_1<b_2$ then $0<b_2-b_1$. If $b_1=b_2$ and $a_1\leq c_1<c_2$ hence $0<c_2-a_1$. If $b_1=b_2$, $c_1=c_2$ and $a_1\leq a_2$ then {$c_2-a_1\geq 0$. Observe that case $c_2-a_1= 0$ then $a_1=b_1=c_1=a_2=b_2=c_2$}. Therefore, if $A\leqot B$ then $B-A\geqot\crispy{0}$. It is not hard to see that the converse is true from Definitions \ref{deftriangularnumero} and \ref{defOT}.
\end{proof}
\begin{observacion}
The order {$\preceq$} proposed in \cite{valentina2021} does not satisfy the previous property {since $(1,2,3)\prec (1,2,4)$, and  $(1,2,4)-(1,2,3)=(-2,0,3)\prec\crispy{0}$}.
\end{observacion}
\vspace{5mm}

\begin{teorema}
Let $A$ and $B$ in $\TFNR$. If $A$ and $B$ are $OT$-positive then $A+B$ is also $OT$-positive.
\end{teorema}

\begin{proof}[Proof.] Let $A=(a_1,b_1,c_1)$ and $B=(a_2,b_2,c_2)$ two $OT$-positive $\TFNR$. Then, from Proposition \ref{proposiciondeB} we have that $b_1\geq 0$, $b_2\geq 0$, $c_1>0$ and $c_2>0$. So, (1) if $b_1>0$ or $b_2>0$ then $b_1+b_2>0$. (2) if $b_1=b_2=0$ then $b_1+b_2=0$ and, because $c_1>0$ and $c_2>0$ and $c_1+c_2>0$. Therefore, in both cases, $A+B>_{OT}\crispy{0}$.
\end{proof}

\begin{teorema}
Let $A$, $B$ and $C$ in $\TFNR$. If $A\leqot B$ then $A+C\leqot B+C$.
\end{teorema}

\begin{proof}[Proof.] Let $A=(a_1,b_1,c_1)$, $B=(a_2,b_2,c_2)$ and $C=(a_3,b_3,c_3)$ be three $\TFN$'s. If  $A\leqot B$ we get the following cases
\begin{enumerate}
\item $b_1<b_2$ or
\item $b_1=b_2$ and $c_1<c_2$ or
\item $b_1=b_2$ and $c_1=c_2$ and $a_1\leq a_2$.
\end{enumerate}
For the first case, we have: If $b_1<b_2$ then $b_1+b_3<b_2+b_3$. Hence $A+C\leqot B+C$. The other cases are analogous.

In this way the Proposition is prove.
\end{proof}
\begin{corolario}\label{sumacontinua}
Let $A$, $B$ $C$ and $D$ be $\TFN$. If $A\leqot B$ and $C\leqot D$ then $A+C\leqot B+D$.
\end{corolario}

{
\begin{proposicion}
Let $A$ and $B$ be two $\TFN$ and $\crispy{r}\in\crispy{\con{R}}$. \begin{enumerate}
\item If $A\leqot B$ and $\crispy{r}\geqot \crispy{0}$ then $\crispy{r}A\leqot \crispy{r}B$,
\item If $A\leqot B$ and $\crispy{r}\leqot \crispy{0}$ then $\crispy{r}A\geqot \crispy{r}B$.
\end{enumerate} 
\end{proposicion}
\begin{proof}
 Straightforward.
\end{proof}
}

\begin{definicion}
{Given $A$, $B$ in $\TFNR$ we} define:
\begin{enumerate}
\item Maximum: \begin{center}
$\Maxx_{\leqot }\{A,B\}=\begin{cases} A,\textrm{ if }B\leqot A;\\
\textrm{ or }\\
B,\textrm{ if }A\leqot B.\end{cases}$
\end{center}
\item Minimum: \begin{center}
$\Minn_{\leqot}\{A,B\}=\begin{cases} B,\textrm{ if }B\leqot A;\\
\textrm{ or }\\
A,\textrm{ if }A\leqot B.\end{cases}
$\end{center}
\end{enumerate}
\end{definicion}

\begin{observacion}
The algebras $(\NFR,+)$ and $(\NFR,\cdot)$,  are not invertible commutative monoids (see \cite[pag.7]{kandasamy2002smarandache}). On the other hand $(\NFR,+,\cdot)$ is not a semiring in the sense \cite{kandasamy2002smarandache}.
\end{observacion}

\section{Triangular Fuzzy Space}\label{TriangularfuzzySpace}
An $n$-tuple of triangular fuzzy numbers $(A_{1},A_{2},A_{3},\ldots,A_{n})$ for an integer $n\geq 1$ is called an $n$-dimensional point or an $n$-dimensional vector. In this case, the $\TFN$'s $A_{1},A_{2},A_{3},\ldots,A_{n}$ are called of coordinates or components of the vector. The collection of all $n$-dimensional vectors is called the vector space of $n$-tuples, or simply $n$-space. We denote this space by $\TFNR^n$. In this paper we shall usually denote vectors by  $\Vector{A},\Vector{B},\Vector{C},\ldots$ and components by the corresponding letters $A$, $B$, $C,~\ldots$ with a subindex in $\{1,...,n\}$. Thus, we write $$\Vector{A}=(A_{1},A_{2},A_{3},\ldots,A_{n}).$$ To convert $\TFNR^n$, into an algebraic system, we introduce two vector operations called addition and multiplication by scalars. The word scalar is used here as a synonym for crisp fuzzy number.

\begin{definicion}\label{operacionesvectoriales}
Two vectors $\Vector{A}$ and $\Vector{B}$ in $\TFNR^n$. The sum $\Vector{A}+\Vector{B}$ is defined to be the vector obtained by adding corresponding components: $$\Vector{A}+\Vector{B}=(A_{1}+B_{1},A_{2}+B_{2},\ldots,A_{n}+B_{n}).$$ If $\crispy{r}\in\crispy{\con{R}}$, we define $\crispy{r}\Vector{A}$ or $\Vector{A}\crispy{r}$ to be the vector obtained by multiplying each component
of $\Vector{A}$ by $\crispy{r}$: $$\crispy{r}\Vector{A}=(\crispy{r}A_{1},\crispy{r}A_{2},\crispy{r}A_{3},\ldots,\crispy{r}A_{n})$$
\end{definicion}
From this definition it is easy to verify the following properties of these operations

\begin{proposicion}\label{operacionesenspacevector}
Vector addition is commutative,
$$\Vector{A}+\Vector{B}=\Vector{B}+\Vector{A},$$
and associative
$$\Vector{A}+(\Vector{B}+\Vector{C})=(\Vector{A}+\Vector{B})+\Vector{C}.$$
Multiplication by scalars is associative,
$\crispy{r}({\crispy{t}\Vector{A}}) = (\crispy{r}\crispy{t})\Vector{A}$
and satisfies the two distributive laws
$$\crispy{r}(\Vector{A}+\Vector{B})={\crispy{r}\Vector{A}+\crispy{r}\Vector{B},}$$ and $$(\crispy{r} + \crispy{t})\Vector{A} ={ \crispy{r}\Vector{A} + \crispy{t}\Vector{A}.}$$
\end{proposicion}

\begin{proof} We will only prove the commutative and distributive law. Let $\Vector{A}$ and $\Vector{B}$ two $\TFNR^n$ and $\crispy{r}\in\crispy{\con{R}}$. For the commutativity we have
\begin{align*}
\Vector{A}+\Vector{B} %= & (A_1,A_2,...,A_n)+(B_1,B_2,...,B_n)\\
 = & (A_1+B_1,A_2+B_2,...,A_n+B_n)\\
  = & (B_1+A_1,B_2+A_2,...,B_n+A_n)\\
 % = & (B_1,B_2,...,B_n)+ (A_1,A_2,...,A_n)\\
    = & \Vector{B}+\Vector{A}.
\end{align*}
On the other hand, for the distributive law we obtain
\begin{align*}
(\crispy{r}+\crispy{t})\Vector{A} = & ((\crispy{r}+\crispy{t})A_1,(\crispy{r}+\crispy{t})A_2,...,(\crispy{r}+\crispy{t})A_n)\\
 = & (\crispy{r}A_1+\crispy{t}A_1,\crispy{r}A_2+\crispy{t}A_2,...,\crispy{r}A_n+\crispy{t}A_n) \,\,\mbox{{ by Remark \ref{rem-ult-sec2}}}\\
= & (\crispy{r}A_1,\crispy{r}A_2,...,\crispy{r}A_n)+(\crispy{t}A_1,\crispy{t}A_2,...,\crispy{t}A_n)\\
%  = & (B_1+A_1,B_2+A_2,...,B_n+A_n)\\
  = & \crispy{r}(A_1,A_2,...,A_n)+ \crispy{t}(A_1,A_2,...,A_n)\\
    = & \crispy{r}\Vector{A}+\crispy{t}\Vector{A}.
\end{align*}
The proof of the remaining properties, we use similar arguments to those given above. Concluding our proof.
\end{proof}
%\rot{The $gH$-difference for fuzzy numbers can be defined as follows:} 
%\begin{definicion} Given $u,v \in \NFR$, the $gH$-difference is the fuzzy number $w$, if it exists, such that $$u\circleddash_{g}v=w \Leftrightarrow \left\{ \begin{array}{lcc} i)\, u=v+w \\ or\, ii)\, v=u+(-1)w \\ \end{array} \right.$$ \end{definicion}

\begin{observacion}\label{observaciondeorden}
The vector with all components $\crispy{0}$ is called the zero vector and is denoted by $\Vector{0}$. It has the property that $\Vector{A}+\Vector{0} = \Vector{A}$ for every vector $\Vector{A}$; in other words, $\Vector{0}$ is an identity element for vector addition. The vector $(-\crispy{1})\Vector{A}$ is also denoted by $-\Vector{A}$ and is called the opposite of $\Vector{A}$ and notice that the equation $\Vector{A}+(-\Vector{A}) = \Vector{0}$ is generally not true. For this, it suffices to consider $\Vector{A}\in \TFNR^n$ wet get
$\Vector{A}-\Vector{A}=(A_1-A_1,...,A_n-A_n)$ from Corollary \ref{teoinverso} have $A_i-A_i>_{OT} \crispy{0}$ whenever $A_i$ is not crisp. Note also that $\Vector{A}\crispy{0} =\Vector{0}$ and that $\Vector{A}\crispy{1}=\Vector{A}$ for all $\Vector{A}\in\TFNR^n$. Therefore, $\TFNR^n$, endowed with the addition and scalar product of Definition \ref{operacionesvectoriales}, is not a vector space on the field of crisp fuzzy numbers. Nevertheless, it is a semi-vector space in the sense of Vasantha Kandasamy \cite{kandasamy2002smarandache} and \cite{milfont2021aggregation}.
\end{observacion}

\begin{definicion}
Let $(A_1,A_2,...,A_n)$ and $(B_1,B_2,...,B_n)$ in $\TFNR^n$. We define the relation $\leq_{\OT^n}$ by
$$(A_1,A_2,...,A_n)\leq_{\OT^n}(B_1,B_2,...,B_n)\Longleftrightarrow \left\{\begin{array}{c}A_1\leqot B_1\textrm{ and }\\
A_2\leqot B_2\textrm{ and }\\ \vdots \\A_n\leqot B_n.\end{array}\right.$$
\end{definicion}

\begin{proposicion}
Let $\Vector{A}\in\TFNR^n$. Then $\Vector{A}-\Vector{A}\geq_{OT^n}\Vector{0}$.
\end{proposicion}
\begin{proof}
Straightforward from Remark \ref{observaciondeorden}.
\end{proof}

The set $\TFNR^n_0$ denote the set of elements of the form $(A_1,...,A_n)$ such that $A_i\in\TFNR_0$ for all $i=1,...n$.

\begin{proposicion}
If $\Vector{A}\in\TFNR^n$ then $\Vector{A}-\Vector{A}\in \TFNR^n_0$.
\end{proposicion}
\begin{proof}
Straightforward from Proposition \ref{proposicionrestadacasicero} and Remark \ref{observaciondeorden}.
\end{proof}

\section{Averaging Functions on $\TFN$}\label{Averaging-Functions}

  \begin{definicion}
A function $\operador{E}{\TFNR^n}{\TFNR}$ is $\OT$-increasing if 
 $$E(\Vector{A})\leqot E(\Vector{B})$$ whenever $\Vector{A}\leq_{\OT^n} \Vector{B}$. 
We can analogously define the $\OT$-decreasing function.
 \end{definicion}
 
Let a function $\operador{E}{\TFNR^n}{\TFNR}$. We call $E$ function type average (in short $\OT$-FTA) if it is $\OT$-increascing and, for each $A_1$, ..., $A_n\in\TFNR$,
$$
\Minn_{\leqot}\{A_1,...,A_n\} \leqot E(A_1,...,A_n) \leqot \Maxx_{\leqot} \{A_1,...,A_n\}.$$

 \begin{definicion}\label{idempotency}
 A function $\operador{E}{\TFNR^n}{\TFNR}$ is idempotent if $E(A,...,A)=A$ for every  $A$ in $\TFNR$.
 \end{definicion}
 
 \begin{proposicion}\label{typeaverange}
 Let $E$ be an $\operador{E}{\TFNR^n}{\TFNR}$. If $E$ is $\OT$-increasing then the following statements are equivalent:
 \begin{enumerate}%[labelindent=\parindent, leftmargin=*,label=]
 \item[(i)] $E$ is idempontent;
 \item[(ii)] $E$ is {an OT}-FTA.
 \end{enumerate}
 \end{proposicion}
 \begin{proof} $(i)\Rightarrow (ii)$ Take any $A_1$, ..., $A_n$ in $\TFNR$. We denote by $A_i=\Minn_{\leqot}\{A_k{;}k=1,...,n\}$ and $A_j=\Maxx_{\leqot }\{A_k{;}k=1,...,n\}$. By monotonicity 
$$A_i=E(A_i,...,A_i)\leqot E(A_1,...,A_n) \leqot E(A_j,...,A_j)=A_j.$$
Hence 
$$\Minn_{\leqot}\{A_k;k=1,...,n\}\leqot E(A_1,...,A_n) \leqot \Maxx_{\leqot }\{A_k;k=1,...,n\}.$$
$(ii)\Rightarrow (i)$ Straightforward.
 \end{proof}
 \begin{definicion}
Let $A_i$ on $\TFNR$ for all $i=1,...,n$. 
\begin{enumerate}%[labelindent=\parindent, leftmargin=*,label=]
\item[(i)] Arithmetic mean
$$M_a(A_1,...,A_n)={\left( \suma{n}{i=1}A_i \right ) \div\crispy{n}}.$$

\item[(ii)] Weighted arithmetic mean
$$M_w(A_1,...,A_n)=\suma{n}{i=1}\crispy{a}_iA_i,$$ where $\crispy{a}_i\geqot\crispy{0}$ for all $i=1,...,n$ and $\suma{n}{i=1}\crispy{a}_i=\crispy{1}$.
\end{enumerate}
 \end{definicion}

  \begin{proposicion}
 The function $\operador{M_w}{\TFNR^n}{\TFNR}$ is {an} $\OT$-FTA.
 \end{proposicion}
 \begin{proof} Let $A_i=\Minn_{\leqot}\{A_k{;}k=1,...,n\}$ and $A_j=\Maxx_{\leqot }\{A_k{;}k=1,...,n\}$ be two $\TFNR$. By monotonicity 
$$A_i=A_i\suma{n}{i=1}\crispy{a}_i\leqot M_w(A_1,...,A_n) \leqot M_w(A_j,...,A_j)=A_j\suma{n}{i=1}\crispy{a}_i=A_j.$$
Hence 
$$\Minn_{\leqot}\{A_k;k=1,...,n\}\leqot M_w(A_1,...,A_n) \leqot \Maxx_{\leqot }\{A_k;k=1,...,n\}.$$
Therefore, by Proposition \ref{typeaverange} we get $M_w$ is {an} $\OT$-FTA.
 \end{proof} 
\begin{corolario}
 The function $\operador{M_a}{\TFNR^n}{\TFNR}$ is {an $\OT$}-FTA.
\end{corolario}
 
%
%\bibliography{bibliografiads}

%\label{bibliografia}

\end{document}